# In Search of Cosmic Time: Complete Observables and the Clock Hypothesis

### v1

@@ 2. Complete Observables and Real Reference Frames
This distinction allows us to disambiguate two important details in the definition of a partial observable, which is often not stressed in the relevant literature. Following Bamonti and Gomes (2024), we should understand partial observables to be relational but gauge-variant quantities that are nevertheless associated measuring procedure. We can understand this seeming contradiction in terms of the fact that partial observables are defined relative to an Idealised Reference Frame. In particular, the parameter of flow equation acts as an IRF and, as such, partial observables are relational in the sense that they describe the correlation between a physical variable and the IRF. Consider our expression for the partial observables (1) above. The variables $q_{i}(t)$ are measurable quantities of the theory but they are not measurable independently of a specification of the value of the flow parameter $t$. Furthermore, since “all measurements are comparisons between different physical systems” (Anderson 1967, p.128), $t$ itself represents a physical system, whose dynamics is neglected as a result of approximations. The second relevant remark is that not every pair of physical quantities to which measuring instruments can be associated can play the role of partial observables. Bona fide partial observables must be dynamically coupled to each other, in order for their relation to constitute a bona-fide complete observable. (Bamonti and Gomes 2024).
@@
Let us then consider the status of the complete observables. In this context, the partial observable that is chosen as the clock observable is playing the role of a reference frame. The point above suggests that to construct a complete observable, we must use DRFs or RRFs. Since we are considering finite dimensional particle mechanics there is no stress-energy tensor to consider. However, the distinction between Dynamical Reference Frames and Real Reference Frames can still be made. That is, a clock variable is always a DRF since its dynamics are always relevant via the flow equation. However, it is only an RRF when the coupling between the clock variable and the other non-clock partial observables is included. Our simple system with free particles is thus an implementation the complete observables programme in terms of a DRF rather than an RRF. [^9] This connection also points to the sense in which the idea of inertial reference frames as discussed in the late nineteenth century by Lange, Neumann, Tait and others are examples of a DRF and not an RRF. This is one way of thinking about Mach’s criticisms. See Thébault (2021a). However, in cases where the coupling is included, complete observables admit an interpretation in terms of an RRF. This is precisely the application of the complete observables approach that we will consider in the context of cosmology in the following section.
@@ 3. Time and the Friedmann Equations
The universe is estimated to be 13.7 billion years old. This estimation is made based upon the standard model of cosmology in which the spacetime structure of the universe is described via general relativity. The base-model of modern cosmology, upon which more sophisticated models are built, is one in which the spatial structure is extremely simple. In particular, following the Cosmological Principle, we assume the universe is homogeneous and isotropic on large scales ($\geq 10^{2}MPc$). This means that, from any location, the distribution of matter and energy appears the same and without any preferred direction. Formally, this corresponds to use of the Friedman (1922)-Lemaître (1931)-Robertson (1935)-Walker (1937) (FLRW) metric, which describes the geometry of spacetime under these symmetries and forms the basis for modelling the large-scale structure and dynamics of the universe.
In the context of the FLRW metric the Einstein Field Equations take on the remarkably simple form given by the Friedmann equations. These equations describe the evolution of a single geometric variable, the scale factor $a(t)$, and take the form:
@@
where $t$ is the so-called cosmological time, $\rho(t)$ and $p(t)$ are the density and pressure of the matter, $k=0,-1,+1$ is the curvature spatial parameter, and the constants have their usual meaning. [^10] The value of $k$ does not fix the overall topology. In fact, different topological choices are possible for the same $k$: for example, a hyperplane (closed topology) is characterised by curvature parameter $k=0$, like a hyperplane (open topology). For a given specification of matter we can solve these equations to get dynamical expressions for the scale factor. For the matter mix that we take to correspond to our universe (including dark energy and dark matter) the relevant expressions describe an expanding universe which matches our observational data to a remarkable degree. [^11] Specifically, observations of the Cosmic Microwave Background (CMB), a relic radiation from the early Universe, provided strong support for the model’s predictions of a hot, dense origin characterising the so-called $\Lambda CDM$ Cosmological Model. Data from missions like COBE, WMAP, and Planck confirmed the homogeneity and isotropy of the CMB within the order of one part in $10^{5}$, in line with the Cosmological Principle. Observations of large-scale structure, such as those from the Sloan Digital Sky Survey (SDSS), further corroborate the predictions made by FLRW cosmology, particularly regarding the distribution of galaxies. These experimental validations solidify the FLRW model as the cornerstone of our understanding of cosmology. This model provides a standard, textbook level story of the expansion of the universe that is assumed by almost all scientists to be unproblematic, at least back to the (presumed) inflationary epoch.
@@
Remarkably, however, when the story regarding the Friedman equations and the expansion of the universe is combined with the Dirac criterion for observables we run into an immediate and deeply problematic conflict. As just noted, the Friedmann equations describe the evolution of the scale factor $a$ and this appears to provide a clear description of the time evolution of spatial geometric structure of the universe. However, the Friedmann equations are equivalent to those generated by a Hamiltonian constraint. So, if the evolution equations generated by a constraint are interpreted as gauge transformations, then we should be understood the Friedmann equations not as dynamical equations by as gauge equations. [^12] We might, of course, simply reject Dirac’s argument connecting gauge transformations to Hamiltonian constraints. In particular, his theorem that first class constraints generate gauge transformations does not apply to Hamiltonian constraints, see (Barbour and Foster 2008) and (Gryb and Thébault 2023, S7.3) for details. Moreover, rigorous formal analysis of these constraints indicate that there are distinct gauge generating and dynamics generating roles that can be explicitly disentangled in the case of theories with a single Hamiltonian constraint. See (Gryb and Thébault 2023, §13) for details. Quantities such as $a$ might appear to evolve over time in cosmology. However, they are not gauge-invariant, and so this evolution is not to be understood as physical. Remarkably, this conflict between the Friedman equations and the Dirac criterion for observables has received almost not detailed discussion in the physics or philosophy literature.
The major exception is the discussion of Thiemann (2006), who explicitly argues that ‘it is incorrect to interpret the FLRW equations as evolution equations of observable quantities’ (p. 9) although he does ‘not doubt the validity the Einstein equations’ he wants to ‘stress that their interpretation as physical evolution equations of observables is fundamentally wrong’ (p. 9). Moreover on this view:
All textbooks on classical GR incorrectly describe the Friedmann equations as physical evolution equations rather than what they really are, namely gauge transformation equations. The true evolution equations acquire possibly observable modifications to the gauge transformation equations whose magnitude depends on the physical clock that one uses to deparametrise the gauge transformation equations.’ (p. 3)
@@
The natural response to this problem is apply the complete and partial observable scheme to construct Dirac observables base upon the Friedmann system of equations. This is precisely what Thiemann suggests to do. His explicit proposal involves introducing a scalar field as a clock, capable of deparametrizing the theory through the Brown and Kuchař (1995)’s mechanism. This approach allows for the construction of a physical Hamiltonian, which generates the evolution of gauge-invariant Dirac observables. Since the dynamics of the chosen material system and its stress-energy tensor are taken into account, it is also to explicitly implement a deparametrization in terms of a Real Reference Frames (RRF) as per the discussion of the last section.
Let us introduce a spatially homogeneous, scalar field $\phi$, which acts as a ‘phantom’ field, not directly observable but capable of generating a physical notion of time. [^13] It is worth nothing here that there are two importantly different senses of phantom that coincide in Thiemann’s usage. First, ‘phantom’ in the sense of ‘missing physics’ that is not directly observable. Second, ‘phantom’ in the more formal sense used by cosmologists as indicating a field with a first order kinetic term in the Lagrangian with a coefficient which has a sign opposite to the sign in the Klein–Gordon Lagrangian. See (Thiemann 2006, p. 4). The key innovation of this approach is to deparametrize the Hamiltonian constraint of GR, transforming it from a constraint equation into a physical Hamiltonian. The Hamiltonian constraint $\mathcal{H}$ is rewritten as: $\mathcal{H}=\pi+\mathfrak{h}$, where where $\pi$ is the conjugate momentum to the scalar field and $\mathfrak{h}$ is called the physical Hamiltonian generating the temporal physical evolution of observables. Using this scalar field Thiemann constructs a framework in which the universe’s time evolution is generated by the physical Hamiltonian rather than the Hamiltonian constraint. This reformulation allows for the deparametrization of the theory, with $\mathfrak{h}$ now acting as a physical Hamiltonian that generates time evolution for gauge-invariant observables. The crucial difference here is that $\mathfrak{h}$ is not constrained to vanish, as is the case with the traditional Hamiltonian constraint in GR.
Once the theory is deparametrized, the time evolution of observables such as the scale factor $a(\phi)$ can be computed using the physical Hamiltonian:
$\frac{da(\phi)}{d\phi}|_{\phi=s}=\{\mathfrak{h},a(\phi)\}|_{\phi=s}\neq 0$
@@
$\{\mathcal{H},a(\phi)\}|_{\phi=s}=0.$
This evolution is now consistent with the Dirac criterion, as $O_{a}(\phi)$ is a Dirac observable that Poisson commutes with all constraints, unlike the original scale factor $a(t)$, which did not Poisson commute with the Hamiltonian constraint.
Crucially, however, in this framework, the evolution of $O_{a}(\phi)$ is governed by a modified version of the Friedmann equations, which includes additional terms due to the presence of the scalar field. In particular, the first Friedmann equations reads as:
$\left(\frac{da/d\phi}{a(\phi)}\right)^{2}=\left[\cfrac{8\pi G}{3}\left[\rho_{m}(\phi)+\rho_{\text{phantom}}(\phi)\right]+\cfrac{\Lambda}{3}\right]\left(1+\frac{1}{x}\right),$
where
$x=\frac{E^{2}}{\alpha^{2}a(\phi)^{6}}$
is a deviation parameter, used to quantify how much the dynamics of the universe, governed by the modified Friedmann equation, differs from the standard cosmological model; $E$ is a constant of motion, representing the energy of the universe; $\alpha$ is a model parameter characterising the influence of the phantom field. We chose $k=0$ to adhere to Thiemann’s formalism.
We thus arrive at an observationally distinct formulation of the theory which implements the Dirac observable prescription. The implication is then that we can either have the standard Friedmann equation and give up on our formalism for gauge invariant observables or we can keep our formalism for observables gauge invariant observables and modify the Friedmann equations. We cannot have both. Thiemann emphasises the gravity of this problem, stating that either the mathematical formalism of GR is inappropriate for cosmology, or we are missing some new physics. In the following section we seek to extricate ourselves from Thiemann’s dilemma based upon the use of Einstein’s famous clock hypothesis: that physical clocks measure proper time along their world-lines.
@@ 4. Hubble Flow and the Clock Hypothesis
Let us return to the derivation of the Friedman equations and seek an alternative physical interpretation of the $t$ in the equations. One crucial aspect of the model we have not yet explicitly considered is the idea of Hubble flow. This flow describes the large-scale motion of matter, driven by the expansion of spacetime itself. More formally, Hubble flow is the component of recessional velocity of matter due to the expansion, separating it from peculiar velocities caused by local gravitational interactions.
@@
One way of understanding the derivation of the FLRW metric is via the adoption of the so-called synchronous frame (Landau and Lifshitz 1987). In this frame there is a common, global cosmological time for all observers comoving with the Hubble flow. Crucially, this means that the synchronous frame is a geodetic reference frame. This means that time-like trajectories, orthogonal to space-like 3D hypersurfaces are geodesics of space-time and the four-velocity of each observer $U^{\mu}=(1,\vec{0})$ automatically satisfy the geodesic equation.
@@
From this perspective the Friedmann equations are not gauge transformation equations. Rather, they are evolution equations in a specific gauge. We can see this as follows. Recall that in the Arnowitt-Deser-Misner (ADM) formalism (Arnowitt et al. 1960) for canonical general relativity the metric is expressed in terms of the lapse function $N(t)$, the shift vector $N^{i}(t)$, and a spatial metric $h_{ij}$. [^14] From Geroch (1970)’s Theorem follows that a globally hyperbolic spacetime can be foliated, that is decomposed into spatial slices parameterized by a global parameter $t$. The ADM formalism can thus be applied to any globally hyperbolic spacetime. The lapse function $N$ formalises the temporal separation between two infinitesimally close hypersurfaces, measured in the normal direction to the first hypersurface. The shift vector $N^{i}$ measures the displacement between the coordinates $x^{i}$ of a point $P\in\Sigma_{t}$ and its orthogonal projection $Q\in\Sigma_{t+dt}$. The connection between $N$ and temporal diffeomorphisms and $N^{i}$ and 3-diffeomorphisms emerges. It is specified that, in order to have a future-directed foliation, the lapse function $N$ must be positive. In this formalism the general line element becomes:
$ds^{2}=-N^{2}dt^{2}+h_{ij}\left(dx^{i}+N^{i}dt\right)\left(dx^{j}+N^{j}dt\right).$
@@
In the specific case of FLRW cosmology, the lapse function $N(t)=1$ and the shift vector $N^{i}(t)=0$. These choices defines the synchronous gauge, reflecting the absence of preferred locations and directions and simplifying the FLRW metric. Since in the synchronous gauge we can assume the coordinate time $t$ to coincide with the proper time $\tau$ measured by observers comoving with the Hubble flow, we will have that the $t$ in the equations will coincide with the proper time of the relevant bundle of geodesics following the Hubble flow. This means we can re-write the equations in terms of proper time $\tau$ simply by equating $t=\tau$. It is worth noting that it is not necessary to show that $\{\mathcal{H},a(\tau)\}\neq 0$, since we already gauge-fixed (deparametrised). So $a(\tau)$ can be seen as a gauge-fixed observable. A gauge-fixed observable is defined in that particular gauge, and need not commute with the constraints because there are no more gauge transformations to refer to. In other words, the gauge freedom has already been eliminated and we no longer have any gauge constraints left, because we have chosen a specific reference frame. To clarifiy what (Dittrich 2007, p.1914) claims, complete observables and gauge-fixed observables are the same, since a choice of reference frame can be seen as a gauge choice (a choice of section on a principle fibre bundle). Or, at least, it is always true that a gauge-fixed observable is a complete observable and only the reverse is less immediately obvious. A more detailed discussion of the status of gauge fixings, reference frames and gauge invariant observables in the context of cosmic proper time, within the fibre bundle formalism, is provided in Appendix A. The important point for our present purposes is to consider the physical consequences of this gauge fixed perspective on the Friedmann equations.
@@
Consider, first that the Hubble flow should be understood to be the flow of a perfect fluid whose stress-energy tensor in the synchronous gauge is $T_{\mu\nu}=\text{diag}(\rho(\tau),-p(\tau))$, where $\rho(\tau)$ is the fluid’s energy density, $p(\tau)$ its pressure. The Friedmann equations are then understood to describe how the expansion rate changes with the proper time of the observers comoving with the fluid. They can be written in terms of the Hubble parameter $H(\tau)=\dot{a}(\tau)/a(\tau)$. The first Friedmann equation is then:
$H(\tau)^{2}=\frac{8\pi G}{3}\rho(\tau)-\frac{kc^{2}}{a(\tau)^{2}}+\frac{\Lambda c^{2}}{3}.$
and the second is:
$\dot{H}(\tau)+H(\tau)^{2}=-\frac{4\pi G}{3}\left(\rho(\tau)+\frac{3p(\tau)}{c^{2}}\right)+\frac{\Lambda c^{2}}{3}$
where we have used the fact that $\frac{\ddot{a}}{a}=\dot{H}+H^{2}$ and assumed differentiation with respect to proper time.

### v2

@@ 2. Complete Observables and Real Reference Frames
A further important disambiguation can be made based upon the connection between partial and complete observables and reference frames. The role of reference frames in general relativity has an extensive philosophical literature. [^11] See Earman and Friedman (1973); Earman (1974), Norton (1989, 1993), DiSalle (2020). Recently, a community of scholars has also emerged in the field of the so-called quantum reference frames, see Giacomini (2021); Kabel et al. (2024) and reference therein. Most relevant to our analysis is the distinction made by Bamonti (2023) between: ‘Idealised Reference Frames’ (IRFs), in which any dynamical interaction of the material system represented by the reference frame is ignored; ‘Dynamical Reference Frames’ (DRFs), in which the set of equations that determine the dynamics of the matter field is included but the the stress-energy tensor of the matter field used as reference frame is neglected; and ‘Real Reference Frames’ (RRFs) in which both the dynamics of the chosen material system and its stress-energy tensor are taken into account. This distinction allows us to disambiguate two important details in the definition of a partial observable, which is often not stressed in the relevant literature. Following Bamonti and Gomes (2024), we should understand partial observables to be relational but gauge-variant quantities that are nevertheless associated with a measuring procedure. We can understand this seeming contradiction in terms of the fact that partial observables are defined relative to an Idealised Reference Frame. In particular, the parameter of flow equation acts as an IRF and, as such, partial observables are relational in the sense that they describe the correlation between a physical variable and the IRF. Consider our expression for the partial observables (1) above. The variables $q_{i}(t)$ are measurable quantities of the theory but they are not measurable independently of a specification of the value of the flow parameter $t$. Furthermore, since “all measurements are comparisons between different physical systems” (Anderson 1967, p.128), [^12] See also (Rovelli 2014, p.99): “In physics, when we talk about measurement, we refer to an interaction between a measured system $S$ and a measuring apparatus $O$.” $t$ itself represents a physical system, whose dynamics is neglected as a result of approximations. The second relevant remark is that not every pair of physical quantities to which measuring instruments can be associated can play the role of partial observables. Bona fide partial observables must be dynamically coupled to each other, in order for their relation to constitute a bona fide complete observable (Bamonti and Gomes 2024).
@@
Let us then consider the status of the complete observables. In this context, the partial observable that is chosen as the clock observable is playing the role of a reference frame. The point above suggests that to construct a complete observable, we must use DRFs or RRFs. Since we are considering finite dimensional particle mechanics there is no stress-energy tensor to consider. However, the distinction between Dynamical Reference Frames and Real Reference Frames can still be made. That is, a clock variable is always a DRF since its dynamics is always relevant via the flow equation. However, it is only an RRF when the coupling between the clock variable and the other non-clock partial observables is included. Our simple system with free particles is thus an implementation of the complete observables programme in terms of a DRF rather than an RRF. [^13] This connection also points to the sense in which the idea of inertial reference frames as discussed in the late nineteenth century by Lange, Neumann, Tait and others are examples of DRFs and not RRFs (see Barbour 2000, pp.101-104 and Barbour 2001, §12). This is one way of thinking about Mach’s criticisms (see Thébault 2021a). However, in cases where the coupling is included, complete observables admit an interpretation in terms of an RRF. This is precisely the application of the complete observables approach that we will consider in the context of cosmology in the following section.
@@ 3. Time and the Friedmann Equations
The universe is estimated to be 13.7 billion years old. This estimation is made based upon the standard model of cosmology — the so-called $\Lambda CDM$ model — in which the spacetime structure of the universe is described via general relativity with a cosmological constant $\Lambda$ (Weinberg 1972). [^14] The ‘CDM’ part stands for ‘Cold Dark Matter’, which is a form of matter that does not interact with electromagnetic radiation (hence ‘dark’) and moves slowly compared to the speed of light (hence ‘cold’). Cold dark matter is not universally regarded as being composed of material particles within an ‘extended Standard Model’. For a detailed discussion, see e.g. Turner (2000). The base-model of modern cosmology, upon which more sophisticated models are built, is one in which the spatial structure is extremely simple. The Cosmological Principle is defined by the condition that the universe is spatially homogeneous and isotropic on large scales ($\geq 10^{2}MPc$). This means that, from any location, the distribution of matter and energy appears the same and without any preferred direction. FLRW spacetimes are the class of generally relativistic spacetimes that are spatially isotropic and homogenous and satisfy certain physically motivated energy conditions or equations of state. [^15] The original papers are Friedman (1922) Lemaître (1931), Robertson (1935), Walker (1937). A formally precise overview is given in (Malament 2012a, §2.11). FLRW spacetimes are the formal basis for modelling the large-scale structure and dynamics of the universe with realistic models involving perturbations about an FLRW metric.
In the context of the FLRW metric and a perfect fluid model of matter, the Einstein Field Equations take on the remarkably simple form given by the Friedmann equations. These equations describe the evolution of a single geometric variable, the scale factor $a(t)$, and take the form:
@@
where $t$ is the cosmological time, $\rho(t)$ and $p(t)$ are the density and pressure of the matter, $k=0,-1,+1$ is the spatial curvature parameter, and the constants have their usual meaning. [^16] The value of $k$ does not fix the overall topology. In fact, different topological choices are possible for the same $k$: for example, a hyperplane (closed topology) is characterised by curvature parameter $k=0$, like a hyperplane (open topology). For a given specification of matter-energy we can solve these equations to get dynamical expressions for the scale factor. For the matter-energy mix that we take to correspond to our universe (including dark energy and dark matter) the relevant expressions describe an expanding universe which matches our observational data to a remarkable degree – although there are existing challenges (Smeenk 2022; Perivolaropoulos and Skara 2022; Colgáin et al. 2024). This model provides a standard, textbook level story of the expansion of the universe that is assumed by almost all cosmologists to be unproblematic, at least back to the (presumed) inflationary epoch.
@@
Remarkably, however, when the story regarding the Friedman equations and the expansion of the universe is combined with the Dirac criterion for observables we run into an immediate and deeply problematic conflict. As just noted, the Friedmann equations describe the evolution of the scale factor $a$ and this appears to provide a clear description of the time evolution of the spatial geometric structure of the universe. However, the Friedmann equations are equivalent to those generated by a Hamiltonian constraint. So, if the evolution equations generated by a constraint are interpreted as gauge transformations, then we should understand the Friedmann equations not as dynamical equations, but as gauge equations. [^17] We might, of course, simply reject Dirac’s argument connecting gauge transformations to Hamiltonian constraints. In particular, his theorem that first class constraints generate gauge transformations does not apply to Hamiltonian constraints, see (Barbour and Foster 2008) and (Gryb and Thébault 2023, §7.3) for details. Moreover, rigorous formal analysis of these constraints indicate that there are distinct gauge generating and dynamics generating roles that can be explicitly disentangled in the case of theories with a single Hamiltonian constraint. See (Gryb and Thébault 2023, §13) for details. Quantities such as $a$ might appear to evolve over time in cosmology. However, they are not gauge-invariant, and so this evolution is not to be understood as physical. Remarkably, this conflict between the Friedman equations and the Dirac criterion for observables has received almost not detailed discussion in the physics or philosophy literature.
@@
We give a sketch of the construction. Let us introduce a spatially homogeneous scalar field $\phi$, which acts as a ‘phantom’ field, which is not directly observable in modern cosmology but can have significant dynamical consequences leading to observable effects. [^18] It is worth nothing here that there are two importantly different senses of phantom that coincide in Thiemann’s usage. First, ‘phantom’ in the sense of ‘missing physics’ that is not directly observable in modern cosmology. Second, ‘phantom’ in the more formal sense used by cosmologists as indicating a field with a first order kinetic term in the Lagrangian with a coefficient which has a sign opposite to the sign in the Klein–Gordon Lagrangian. See (Thiemann 2006, p. 4). The key innovation of this approach is to deparametrise the Hamiltonian constraint of GR, transforming it from a constraint equation into a physical Hamiltonian. The Hamiltonian constraint $\mathcal{H}$ is rewritten as: $\mathcal{H}=\pi+\mathtt{h}$, where $\pi$ is the conjugate momentum to the scalar field and $\mathtt{h}$ is called the physical Hamiltonian generating the temporal physical evolution of observables. Using this scalar field Thiemann constructs a framework in which the universe’s time evolution is generated by the physical Hamiltonian rather than the Hamiltonian constraint. This reformulation allows for the deparametrisation of the theory, with $\mathtt{h}$ now acting as a physical Hamiltonian that generates time evolution for gauge-invariant observables. The crucial difference here is that $\mathtt{h}$ is not constrained to vanish, as is the case with the traditional Hamiltonian constraint in GR.
Once the theory is deparametrised, the time evolution of observables such as the scale factor $a(\phi)$ can be computed using the physical Hamiltonian:
(7) $\frac{da(\phi)}{d\phi}|_{\phi=s}=\{\mathtt{h},a(\phi)\}|_{\phi=s}\neq 0$
@@
(8) $\{\mathcal{H},a(\phi)\}|_{\phi=s}=0.$
This evolution is now consistent with the Dirac criterion, as $a(\phi)$ is a Dirac observable that Poisson commutes with all constraints, unlike the original scale factor $a(t)$, which did not Poisson commute with the Hamiltonian constraint. Crucially, however, in this framework, the evolution of $a(\phi)$ is governed by a modified version of the Friedmann equations, which includes additional terms due to the presence of the scalar field. In particular, the first Friedmann equations reads as:
(9) $\left(\frac{da/d\phi}{a(\phi)}\right)^{2}=\left[\cfrac{8\pi G}{3}\left[\rho_{m}(\phi)+\rho_{\text{phantom}}(\phi)\right]+\cfrac{\Lambda}{3}\right]\left(1+\frac{1}{x}\right),$
where $x=\frac{E^{2}}{\alpha^{2}a(\phi)^{6}}$ is a deviation parameter, used to quantify how much the dynamics of the universe, governed by the modified Friedmann equation, differs from the standard cosmological model; $E$ is a constant of motion, representing the energy of the universe; $\alpha$ is a model parameter characterising the influence of the phantom field. We chose $k=0$ to adhere to Thiemann’s formalism.
We thus arrive at an observationally distinct formulation of the theory which implements the Dirac observable prescription. The implication is then that we can either have the standard Friedmann equation and give up on our formalism for gauge-invariant observables or we can keep our formalism for observables gauge-invariant observables and modify the Friedmann equations. We cannot have both. Thiemann emphasises the gravity of this problem, stating that either the mathematical formalism of GR is inappropriate for cosmology, or we are missing some new physics. In the following section we seek to extricate ourselves from Thiemann’s dilemma based upon the use of Einstein’s famous clock hypothesis: that physical clocks measure proper time along their world-lines.
@@ 4. Hubble Flow and the Clock Hypothesis
Let us return to the derivation of the Friedman equations and seek an alternative physical interpretation of the $t$ in the equations. One crucial aspect of the model we have not yet explicitly considered is the idea of Hubble flow. This ‘flow’ describes the large-scale motion of matter, driven by the expansion of spacetime itself. More formally, Hubble flow is the component of recessional velocity of matter due to the expansion, separating it from peculiar velocities caused by local gravitational interactions.
@@
One way of understanding the derivation of the FLRW metric is via the adoption of the so-called synchronous reference frame (Landau and Lifshitz 1987). In this frame there is a common, global cosmological time for all observers comoving with the Hubble flow. Crucially, this means that the synchronous frame is a geodetic reference frame. Consequently, time-like trajectories orthogonal to space-like 3D hypersurfaces are geodesics of space-time and the four-velocity of each observer $U^{\mu}=(1,\vec{0})$ automatically satisfy the geodesic equation.
@@
From this perspective the Friedmann equations are not gauge transformation equations. Rather, they are evolution equations in a specific gauge. We can see this as follows. Recall that in the Arnowitt-Deser-Misner (ADM) formalism (Arnowitt et al. 1960) for canonical general relativity the metric is expressed in terms of the lapse function $N(t)$, the shift vector $N^{i}(t)$, and a spatial metric $h_{ij}$. [^19] From Geroch (1970)’s Theorem follows that a globally hyperbolic spacetime can be foliated, that is decomposed into spatial slices parametrised by a global parameter $t$. The ADM formalism can thus be applied to any globally hyperbolic spacetime. The lapse function $N$ formalises the temporal separation between two infinitesimally close hypersurfaces, measured in the normal direction to the first hypersurface. The shift vector $N^{i}$ measures the displacement between the spatial coordinates $x^{i}$ of a point $P\in\Sigma_{t}$ and its orthogonal projection $Q\in\Sigma_{t+dt}$. The connection between $N$ and temporal diffeomorphisms and $N^{i}$ and 3-diffeomorphisms emerges. It is specified that, in order to have a future-directed foliation, the lapse function $N$ must be positive. In this formalism the general line element becomes:
(10) $ds^{2}=-N^{2}dt^{2}+h_{ij}\left(dx^{i}+N^{i}dt\right)\left(dx^{j}+N^{j}dt\right).$
In the specific case of FLRW cosmology, the lapse function $N(t)=1$ and the shift vector $N^{i}(t)=0$.
These choices defines the synchronous gauge, reflecting the absence of preferred locations and directions and simplifying the FLRW metric. Since in the synchronous gauge we can assume the coordinate time $t$ to coincide with the proper time $\tau$ measured by observers comoving with the Hubble flow, we will have that the $t$ in the equations will coincide with the proper time of the relevant bundle of geodesics following the Hubble flow. This means we can re-write the equations in terms of proper time $\tau$ simply by equating $t=\tau$.
@@
It is worth noting that it is not necessary to show that $\{\mathcal{H},a(\tau)\}=0$, since we already ‘gauge-fixed’ to the synchronous gauge. This means that $a(\tau)$ can be seen as a gauge-fixed observable and, as such, is gauge-invariant. A gauge-fixed observable is defined in that particular gauge, and need not commute with the constraints because there are no more gauge transformations to refer to. In other words, the gauge freedom has already been eliminated and we no longer have any gauge constraints left, because we have chosen a specific reference frame (on this, see e.g. Wallace 2024). When can then clarify the claim of (Dittrich 2007, p.1914) that complete observables and gauge-fixed observables are the same, since a choice of reference frame can be seen as a gauge choice (a choice of section on a fibre bundle). Or, at least, it is always true that a gauge-fixed observable is a complete observable and only the reverse is less immediately obvious. For our current purpose, the key insight is recognising that the Friedmann equations, when considered in a synchronous gauge, function like dynamical evolution equations. [^20] A more detailed discussion of the status of gauge-fixings, reference frames and gauge invariant observables in the context of cosmic proper time, within the fibre bundle formalism, is provided in Appendix Af. It is worth noting that conceptualising the Friedmann equations as gauge-fixed equations can be made consistent with the arguments of (Gryb and Thébault 2023, §13.2) which imply a dynamical view of global Hamiltonian constraints in which the flow along a solution of the Friedmann equations would not formally be a gauge symmetry. This is because one may view the symmetries of the theory as acting on entire histories and take their action on the boundary to be fixed and then recover the ‘gauge-fixed’ perspective where the Friedmann equations in synchronous gauge describe evolution of the true degrees of freedom.
@@
To show this, consider that the Hubble flow should be understood to be the flow of a perfect fluid whose stress-energy tensor in the synchronous gauge is $T_{\mu\nu}=\text{diag}(\rho(\tau),-p(\tau))$, where $\rho(\tau)$ is the fluid’s energy density, $p(\tau)$ its pressure. The Friedmann equations are then understood to describe how the expansion rate changes with the proper time of the observers comoving with the fluid. They can be written in terms of the Hubble parameter $H(\tau)=\dot{a}(\tau)/a(\tau)$ as:
(11) $H(\tau)^{2}=\frac{8\pi G}{3}\rho(\tau)-\frac{kc^{2}}{a(\tau)^{2}}+\frac{\Lambda c^{2}}{3}$
(12) $\dot{H}(\tau)+H(\tau)^{2}=-\frac{4\pi G}{3}\left(\rho(\tau)+\frac{3p(\tau)}{c^{2}}\right)+\frac{\Lambda c^{2}}{3}$